OpenStudy (anonymous):

Anmol spends 20% of her monthly income on children's education and 15% on food. 40% of the remaining she spends on rent and transportation together and 30% on entertainment. Now she is left with an amount of Rs. 8775 after all these expenditures. Find her yearly income.

OpenStudy (anonymous):

27.3% left so 27.3x/100=8775

OpenStudy (unklerhaukus):

im not sure how you got that/. 100% -20% on education , -15% on food how much percent is remaining at this stage ?

OpenStudy (anonymous):

65%

OpenStudy (unklerhaukus):

yeah thats right , so if 40% of 65% spends on rent and transportation and what percentage of the whole monthly income is this?

OpenStudy (anonymous):

26

OpenStudy (unklerhaukus):

good, and what about entertainment?

OpenStudy (anonymous):

39*30%

OpenStudy (anonymous):

27.3 % left with

OpenStudy (unklerhaukus):

i think the percentage of monthly income spent on entertainment is 65%*30%

OpenStudy (unklerhaukus):

because it is 30% of the remaining 65%

OpenStudy (anonymous):

40%?

OpenStudy (unklerhaukus):

?

OpenStudy (anonymous):

onrent and transport

OpenStudy (unklerhaukus):

after we took (15+20)% away from 100% we had 65% remaining 30% of this is spent on entertainment

OpenStudy (anonymous):

40 % on rent and transport

OpenStudy (hesan):

@dg2 What you have to find monthly income or yearly income

OpenStudy (anonymous):

yearly income

OpenStudy (anonymous):

i got it @UnkleRhaukus

OpenStudy (anonymous):

(100-20+15/100)*(100-40+30/100)x=8775 x=45000

OpenStudy (anonymous):

45000 per month and 45000*12=540000 yearly

OpenStudy (anonymous):

u got it?

OpenStudy (unklerhaukus):

ya thats what i got!